OpenStudy (the_puppet1):

What is the solution of the equation n + 13 = 30? A) n = 13 B) n = 17 C) n = 30 D) n = 43

OpenStudy (cas_fangirl_14):

Its B

OpenStudy (squidnnneyyy):

Ok so for a problem where you solve for a variable, you have to get the variable on one side of the equal sign by itself. To do this you have to do the opposite of what the question asks

OpenStudy (moldybubblegum12):

n = 17 you just have to subtract 13 from 30

OpenStudy (squidnnneyyy):

Since it says n+13, you do the opposite of addition

OpenStudy (the_puppet1):

But guys plz don't have me get it wrong my mom says I need a 120% on extra credit

OpenStudy (squidnnneyyy):

The opposite of addition is subtraction so you subtract 13 from both sides. n+13=30 -13 -13 When you subtract 13 is canceled out so you have n=30-13. 30-13 is 17. n=17